OpenStudy (anonymous):

Simplify. 8^–2 • 8^14 A. 8^12 B. 8^–28 C. 8^–7 D. 8^16

OpenStudy (mathstudent55):

This is the rule of exponents you need for this problem. \(a^m \cdot a^n = a^{m + n} \)

OpenStudy (solomonzelman):

ADD THE EXPONENTS.

OpenStudy (mathstudent55):

Example: \(5^{-3} \cdot 5^{13} = 5^{-3 + 13} = 5^{10} \)

OpenStudy (anonymous):

When you have a multiplication (like the one that's showing), then you add. When you have a division you subtract. Multiply = Add Divide = Subtract

OpenStudy (solomonzelman):

\(\Huge\color{blue}{ \bf a^b \times a^{c}=a^{b+c} }\) your a is 8, your b is -2, your c is 14.

OpenStudy (anonymous):

okay, thank you :)

OpenStudy (mathstudent55):

You're welcome.